When 4.37 g of potassium bromide (KBr) is dissolved in 109 g of water in a styrofoam calorimeter of negligible heat capacity, the temperature drops from 25.00 to 23.34 °C. Based on this observation, calculate q for the water and ΔH° for the process, assuming that the heat absorbed by the salt is negligible.

  • KBr(s) --> K+(aq) + Br- (aq)

The specific heat of water is 4.184 J °C-1 g-1.
Answer #1

4.37 g of potassium bromide (KBr) = mass / molar mass = 4.37 g / 119.002 g/mole = 0.0367 mole.

heat gain by water (q) = m * s * dT = 109 * 4.184 * (23.34 - 25.00) = - 757.1 J

ΔH° for the process = 757.1 J / 0.0367 mole = 20628 J / mole = 20.6 KJ / mole.